A man says truth 3 out of 4 times..he throws a dice nd says its a six..what is the probability it was actually a six..now by bayes theorem if we solve we get 3/8, but by simple logic..he says SOMETHING HAS HAPPE.NED..nd we know 3 out of 4 times SOMETHING actually HAS HAPPENNED..so just like for any other statement..the statement " dice shows 6 " will be true 3/4 times..whats wrng in the reasoning..please help!!!

Grade:12

2 Answers

Aman Bansal
10 years ago

Dear Swapnam,

P(it is six|reports six) = P(it is 6, reports six)/P(reports
six) = [3/4*1/6] / [(3/4*1/6) + (1/4*5/6*1/5)] = 3/4
Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 
http://www.askiitians.com/packages/packages.aspx
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS Thanks Aman Bansal Askiitian Expert
Yogita Bang
10 years ago

Dear Aman sir,

In ur answer

P(it is six|reports six) = P(it is 6, reports six)/P(reports six) = [3/4*1/6] / [(3/4*1/6) + (1/4*5/6*1/5)] = 3/4

From where (underlined) 1/5 comes......... PLEASSE EXPLAIN